OpenStudy (anonymous):

An auditorium has 30 rows of seats. The first row has 20 seats, the second row has 22 seats, the third row has 24 seats, and so on. What is the total seating capacity of the auditorium?

OpenStudy (anonymous):

sum = n (a1 + an)/2, n = number of rows a1 = number of seat in the first row an = number of seat in the 30th row

OpenStudy (kagıtucak):

20+22+24+....... 2(10+11+12+.....)=2(30.31/2-9.10/2)=30.31-9.10=930-90=840

OpenStudy (anonymous):

cant be 840, can only be 1,200 1,470 1,530 1,640

OpenStudy (kagıtucak):

i used 1+2+........+n=n(n+1)/2

OpenStudy (anonymous):